OpenStudy (anonymous):

Can someone help? Find the intersection of the line through the points (7,6) and (4,2) with the line y=x . The point of intersection is (A,B) . What is the value of A and B?

hartnn (hartnn):

can you find the equation of line through the points (7,6) and (4,2) ?

hartnn (hartnn):

first you need to find slope , The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

m=4/3 right?

hartnn (hartnn):

absolutely! now equation of line will be y-y1 = m (x-x1) you can use any of the 2 points for x1,y1.

hartnn (hartnn):

can u find equation of line ? or need help with that ?

OpenStudy (anonymous):

y=4/3x-2 2/3 is that right?

OpenStudy (anonymous):

y-2=4/3(x-4) that is what I had at first.

hartnn (hartnn):

thats correct, to find the intersection point you need to solve both equations simultaneously, that is put x=y in y-2=4/3(x-4) and find the value of x (or y)

hartnn (hartnn):

then x=y=A=B

OpenStudy (anonymous):

A=10 B=10